# Natural Language Processing for the Legal Domain: A Survey of Tasks, Datasets, Models, and Challenges

### v1

@@ 4.5.1. Task
Dozier et al. (2010) present pioneering work in NER within legal documents such as US case law and pleadings, employing three methodologies: lookup, contextual rules, and statistical models to detect entities like judges, attorneys, and legal terms. Their system adapts these approaches to the specialized context of legal texts, processing various types of documents and extracting legal entities. This work highlights the challenges and necessary adaptations for deploying NER in the legal domain, where the specialized language and high accuracy are required for successful legal analysis.
@@
Păis et al. (2021) develop a NER model for the legal domain that integrates Bi-LSTM cells and a Conditional Random Fields (CRFs) layer, utilizing multiple data sources and embedding types. Their system architecture combines word embeddings from pre-trained models, character embeddings, gazetteer resources from the GeoNames [^6] https://www.geonames.org/ database and JRC-Names (Steinberger et al., 2011), and known affixes to enrich the model’s understanding of legal text. Their training process involved fine-tuning the word embeddings, while dynamically learning character embeddings with subsequent Bi-LSTM layers, enhancing the model’s capability to generalize across unseen texts. They implemented the system using a modified version of NeuroNER (Dernoncourt et al., 2017), allowing for online model serving and incorporating features like dropout for regularization and gradient clipping to handle exploding gradients. They also explored ensemble methods to improve model accuracy by combining results from multiple model configurations, measuring the effectiveness through precision, recall, and F1 scores against a gold standard corpus.
@@
Smădu et al. (2022) explore domain adaptation in Legal NER, focusing on the Romanian and German languages. They utilize a model architecture that integrates a pre-trained BERT (Devlin et al., 2019) layer for feature extraction with Bi-LSTM networks to handle sequence dependencies and CRFs for sequence tagging. Their approach employs domain adaptation techniques through a gradient reversal layer connected to a domain discriminator, aimed at reducing domain-specific biases and enhancing feature transferability across domains. This model trains on both legal and general domains simultaneously, adapting to the peculiarities of each through adversarial learning that modifies the learning process dynamically. This setup allows for improved generalization of the NER system across varied linguistic and domain contexts, though the actual performance varied with minimal improvements noted for German and decreased performance for the Romanian legal dataset.
4.5.2. Datasets
@@
Leitner et al. (2020) present a dataset for NER focused on German federal court decisions, containing approximately 67,000 sentences and over two million tokens. This dataset features 54,000 manually annotated entities distributed across 19 fine-grained semantic classes, specifically tailored to the legal domain, such as court, judge, lawyer, law, person, and legal literature, along with over 35,000 TimeML-based time expressions. The annotations cover both broad categories like location, person, and organization, and more specialized ones like legal norms and case-by-case regulations, distinguishing between different types of legal acts and literature. This dataset’s comprehensive annotation process involved multiple cycles to refine the tagging guidelines and enhance annotation quality.
@@
Păis et al. (2021) introduce the LegalNERo corpus, a manually annotated resource for NER in the Romanian legal domain, featuring 370 legal documents annotated with five general entity types: person, location, organization, time expressions, and legal references. This corpus was developed to support both specific legal domain NER tasks and more general NER applications by enabling compatibility with existing general-purpose NER systems. The corpus includes rich entity annotations, with legal references showing the highest token count per entity, indicating their complexity and length. The detailed annotation process, including inter-annotator agreement assessed by Cohen’s Kappa, and the subsequent mapping of entities to RDF format, highlights the corpus’s utility and precision for advancing NER research and applications within the legal domain.
@@
Au et al. (2022) introduce the E-NER dataset, an annotated collection derived from the US SEC’s EDGAR filings, designed for legal NER. This dataset contains filings that are rich in text, such as quarterly reports (Form 10-Q) and significant event announcements (Form 8-K), from which sentences were extracted and annotated with seven named entity classes more tailored to legal content than those in the standard CoNLL dataset(Tjong Kim Sang and De Meulder, 2003). The entities include Person, Location, Organization, Government, Court, Business, and Legislation/Act, adjusting the CoNLL classes to better suit legal documents. E-NER contains significantly longer sentences compared to CoNLL and includes detailed annotations of financial entities from legal company filings.
@@
Kalamkar et al. (2022) present a comprehensive corpus aimed at enhancing legal NER, containing 46,545 entities across 14 types identified in Indian High Court and Supreme Court judgments. This corpus, split into preamble and judgment sections, includes diverse entity types detailed in their legal NER taxonomy, such as court, petitioner, respondent, and statute, among others. The training set, drawn from judgments between 1950 and 2017, features 29,964 entities and the development and test sets, spanning 2018 to 2022. This dataset not only facilitates training and evaluation of NER models specific to the legal domain but also provides a structured framework for assessing the performance of NER systems on legal texts. Their approach leverages a combination of manual annotation and ML techniques to ensure the precision of entity recognition in legal judgments.
4.6. Large Legal Corpora
The foundational step in training an LLMs is the use of extensive corpora. For the development of a sophisticated LLMs that effectively addresses a wide range of legal NLP tasks, it is crucial to have access to large-scale legal corpora. These corpora must meet several critical criteria to ensure their effectiveness and ethical utility. First, they should be transparent in their sourcing and composition, allowing users to understand the origins and types of included data. Additionally, the privacy of individuals should be safeguarded, preventing any potential invasion of personal data. It is also imperative to minimize toxicity and bias within the corpora to promote fairness and accuracy in model outcomes. By following these principles, we can enhance the capabilities of LLMs in the legal domain. This ensures they are both powerful and reliable tools for legal analysis. In this section, we will explore the existing large legal corpora and online databases.
@@ 4.6. Large Legal Corpora
Zheng et al. (2021) introduce the CaseHOLD dataset, a novel benchmark for evaluating NLP models in the legal domain, designed to address the challenge of identifying the legal holdings from case texts. The dataset contains over 53,000 multiple choice questions derived from U.S. case law citations, where each question requires the identification of the correct holding from a set of potential answers. This task, simulating a fundamental lawyering skill taught in law school, involves contextual understanding and application of legal rules to factual situations. CaseHOLD is aimed at enhancing model training by focusing on semantic matching and the ability to discern nuanced legal principles. The dataset is structured to provide a challenging yet accessible resource for NLP researchers, with a clear focus on promoting deeper understanding and application of legal rules in automated systems. It utilizes a format where a cited text serves as a prompt with five answer options—one correct holding and four closely related incorrect holdings—to refine the models’ abilities to accurately reflect legal reasoning.
@@
Chalkidis et al. (2022a) introduce the Legal General Language Understanding Evaluation (LexGLUE) benchmark, a comprehensive suite of datasets aimed at assessing the capabilities of NLP models across various legal tasks. The benchmark covers datasets such as ECtHR (Chalkidis et al., 2019a), SCOTUS [^7] https://www.supremecourt.gov, EUR-LEX, LEDGAR (Tuggener et al., 2020), UNFAIR-ToS (Lippi et al., 2019), and CaseHOLD (Zheng et al., 2021), each chosen for its complexity, relevance, and need for legal expertise. These datasets cover a range of tasks from multi-label and multi-class classification to multiple-choice questions and are split chronologically into training, development, and test sets to provide standardized evaluation metrics. For instance, ECtHR datasets focus on violations of European Convention of Human Rights provisions, SCOTUS database classifies U.S. Supreme Court opinions by legal issues, EUR-LEX database involves labeling EU laws with EuroVoc concepts, LEDGAR classifies provisions of U.S. contracts, UNFAIR-ToS identifies unfair terms in online service agreements, and CaseHOLD involves answering questions about legal rulings. This benchmark facilitates the testing of NLP models, addressing the challenges of legal text comprehension and understanding required for effective application in the legal domain.
@@
Chalkidis et al. (2022b) present FairLex, a benchmark suite consisting of four legal datasets—ECtHR (Chalkidis et al., 2019a), SCOTUS, FSCS, and CAIL (Xiao et al., 2018)—that address the fairness of NLP applications across diverse legal jurisdictions and languages, including English, German, French, Italian, and Chinese. These datasets, curated from European Council, USA, Switzerland, and China, cover various legal tasks such as judgment prediction, issue area classification, and crime severity prediction, aiming to test the performance and fairness of LMs in recognizing and classifying legal texts. FairLex focuses on ensuring demographic, regional, and legal topic fairness by analyzing attributes like gender, age, region of origin, and legal areas within cases. Each dataset in FairLex provides a substantial number of cases, systematically divided into training, development, and test sets, and includes detailed attributes like the defendant state in ECtHR, decision direction in SCOTUS, legal areas in FSCS, and demographic details in CAIL. The CAIL dataset from China contains over a million cases focusing on criminal law, annotated with demographics and regional classifications, which are used to explore the crime severity prediction task.
Henderson et al. (2022) introduce the ‘Pile of Law’, the first and an important large corpus in the legal domain, containing a 256GB dataset of open-source English-language legal and administrative data. This dataset includes contracts, court opinions, legislative records, and administrative rules, curated to explore data sanitization norms across legal and administrative settings and serve as a tool for pre-training legal-domain LMs. They emphasize the legal norms governing privacy and toxicity filtering, detailing how the dataset reflects these norms through built-in filtering mechanisms in the collected data, which include court filings, legal analyses, and government publications. By analyzing how legal and administrative entities handle sensitive information and potentially offensive content, the paper provides actionable insights for researchers to improve content filtering practices before pre-training LLMs, thereby enhancing the ethical use of NLP in legal applications.
@@
Rabelo et al. (2022) summarize the 8th Competition on Legal Information Extraction and Entailment (COLIEE 2021), which featured five tasks across case and statute law, engaging participants from various teams to apply diverse NLP approaches. The competition tasks included case law retrieval and entailment, as well as statute law retrieval and entailment with and without prior retrieved data. Specifically, Task 1 focused on extracting relevant supporting cases from a corpus, while Task 2 involved identifying paragraphs from cases that entail a given new case fragment. For statute law, Tasks 3 and 4 entailed retrieving and answering questions based on civil code statutes, with Task 5 challenging participants to answer without pre-retrieved statutes. The datasets used varied in complexity, from 4415 case files in Task 1 with a need to identify noticed cases without relying on citations, to the civil code-based Tasks 3, 4, and 5 which adapted to recent legal revisions in Japanese law and excluded untranslated parts, reflecting the ongoing evolution and challenge in legal NLP applications.
@@
Barale et al. (2023) present AsyLex, a pioneering dataset tailored for Refugee Law applications, featuring 59,112 documents from Canadian refugee status determinations spanning from 1996 to 2022. This dataset is designed to enhance the capabilities of NLP models in legal research by providing 19,115 gold-standard human-annotated and 30,944 inferred labels for entity extraction and LJP. Key contributions include anonymizing decision documents, employing a robust annotation methodology, and creating datasets for specific NLP tasks like entity extraction and judgment prediction. This rich corpus, with detailed annotations across 22 categories, supports complex legal NLP tasks, thereby filling the gap in resources for the legal domain.
@@
Niklaus et al. (2023) introduce LEXTREME, a multilingual benchmark specifically designed to evaluate LMs on legal NLP tasks, a critical step given the unique challenges of legal language. Surveying legal NLP literature from 2010 to 2022, they curate 11 datasets spanning 24 languages and cover a variety of legal domains, employing datasets that only involve human-annotated texts or those with annotations derived through clear methodological frameworks. They introduce two aggregate scores to facilitate fair comparison across models: the dataset aggregate score and the language aggregate score, revealing a performance correlation with model size on LEXTREME. The benchmark consists of three task types: Single Label Text Classification, Multi Label Text Classification, and NER, using existing splits for training, validation, and testing when available, or creating random splits otherwise. This effort marks a significant advancement in testing NLP capabilities across a diverse range of legal documents and languages.
@@
Park and James (2023) explore the creation of a Natural Language Inference dataset within the legal domain, focusing on criminal court verdicts in Korean. Their methodology includes the innovative use of adversarial hypothesis generation to challenge annotators and enhance the robustness of the dataset, supported by visual tools for hypothesis network construction. The data collection involves extracting context from verdicts and augmenting it using Easy Data Augmentation (Wei and Zou, 2019) techniques and round-trip translation to generate a dataset for training and testing Natural Language Inference models. The study highlights issues such as annotators’ limited domain knowledge and challenges in handling long contexts but provides solutions like targeted data collection and the use of gamification to boost annotator engagement and productivity.
@@
Goebel et al. (2024) summarize the 10th Competition on Legal Information Extraction and Entailment (COLIEE 2023), featuring four tasks across case and statute law with participation from ten different teams engaging in multiple tasks. Task 1 involves legal case retrieval, requiring participants to extract supporting cases from a corpus, and Task 2 focuses on legal case entailment, identifying paragraphs that entail aspects of a new case. Task 3 and 4, based on Japanese Civil Code statutes from the bar exam, involve retrieving relevant articles and verifying statements, respectively. The competition leverages a dataset of over 5,700 case law files and introduces new query cases and test questions sourced from recent bar exams, testing the efficacy of different teams’ approaches in handling complex legal texts and hypotheses in a controlled competitive environment.
@@
Östling et al. (2024) introduce the Cambridge Law Corpus (CLC), a comprehensive legal dataset featuring 258,146 cases from UK courts, dating from the 16th century to the present. The corpus includes raw text and metadata across various court types, and is structured in XML format for ease of use and annotated for case outcomes in a subset of 638 cases. Additionally, the CLC is supported by a Python library for data manipulation and ML applications.

### v2

@@ 9.1. Task
Dozier et al. (2010) present pioneering work in NER within legal documents such as US case law and pleadings, employing three methodologies: lookup, contextual rules, and statistical models to detect entities, such as judges, attorneys, and legal terms. Their system adapts these approaches to the specialised context of legal texts, processing various types of documents and extracting legal entities. This work highlights the challenges and necessary adaptations for deploying NER in the legal domain, where the specialised language and high accuracy are required for successful legal analysis.
@@
Păis et al. (2021) develop a NER model for the legal domain that integrates Bi-LSTM cells and a Conditional Random Fields (CRFs) layer, utilising multiple data sources and embedding types. Their system architecture combines word embeddings from pre-trained models, character embeddings, gazetteer resources from the GeoNames [^7] https://www.geonames.org/ database and JRC-Names (Steinberger et al., 2011), and known affixes to enrich the model’s understanding of legal text. Their training process involved fine-tuning the word embeddings, while dynamically learning character embeddings with subsequent Bi-LSTM layers, enhancing the model’s capability to generalise across unseen texts. They implemented the system using a modified version of NeuroNER (Dernoncourt et al., 2017), allowing for online model serving and incorporating features, such as dropout for regularisation and gradient clipping to handle exploding gradients. They also explored ensemble methods to improve model accuracy by combining results from multiple model configurations, measuring the effectiveness through precision, recall, and F1 scores against a gold standard corpus.
@@
Smădu et al. (2022) explore domain adaptation in legal NER, focusing on the Romanian and German languages. They utilise a model architecture that integrates a pre-trained BERT layer for feature extraction with Bi-LSTM networks to handle sequence dependencies and CRFs for sequence tagging. Their approach employs domain adaptation techniques through a gradient reversal layer connected to a domain discriminator, aimed at reducing domain-specific biases and enhancing feature transferability across domains. This model trains on both legal and general domains, adapting to the peculiarities of each through adversarial learning that modifies the learning process dynamically. This setup allows for improved generalisation of the NER system across varied linguistic and domain contexts, though the actual performance varied with minimal improvements noted for German and decreased performance for the Romanian legal dataset.
Adhikary et al. (2023) publish LeDA, a legal data annotation system designed to address the challenges of extracting legal entities and concepts from case documents. Unlike traditional sequence labelling tools, LeDA enables annotators to dynamically define new legal concepts during annotation. The system also incorporates a meta-annotation mechanism for adjudicating conflicting annotations.
9.2. Datasets
@@
Leitner et al. (2020) present a dataset for NER focused on German federal court decisions, containing approximately 67,000 sentences and over two million tokens. This dataset features 54,000 manually annotated entities distributed across 19 fine-grained semantic classes, specifically tailored to the legal domain, such as court, judge, lawyer, law, person, and legal literature, along with over 35,000 TimeML-based time expressions. The annotations cover both broad categories, such as location, person, and organisation, and more specialised ones, such as legal norms and case-by-case regulations, distinguishing between different types of legal acts and literature. This dataset’s annotation process involved multiple cycles to refine the tagging guidelines and enhance annotation quality.
@@
Păis et al. (2021) introduce the LegalNERo corpus, a manually annotated resource for NER in the Romanian legal domain, featuring 370 legal documents annotated with five general entity types: person, location, organisation, time expressions, and legal references. This corpus was developed to support both specific legal domain NER tasks and more general NER applications by enabling compatibility with existing general-purpose NER systems. The corpus includes rich entity annotations, with legal references showing the highest token count per entity, indicating their complexity and length. The detailed annotation process, including inter-annotator agreement assessed by Cohen’s Kappa, and the subsequent mapping of entities to RDF format, highlights the corpus’s utility and precision for advancing NER research and applications within the legal domain.
@@
Au et al. (2022) introduce the E-NER dataset, an annotated collection derived from the US Securities and Exchange Commission’s EDGAR filings, designed for legal NER. This dataset contains filings that are rich in text, such as quarterly reports and significant event announcements, from which sentences were extracted and annotated with seven named entity classes more tailored to legal content than those in the standard CoNLL dataset (Tjong Kim Sang and De Meulder, 2003). The entities include person, location, organisation, government, court, business, and legislation/act, adjusting the CoNLL classes to better suit legal documents. E-NER contains longer sentences compared to CoNLL and includes detailed annotations of financial entities from legal company filings.
@@
Kalamkar et al. (2022) present a corpus aimed at enhancing legal NER, containing 46,545 entities across 14 types identified in Indian High Court and Supreme Court Judgements. This corpus, split into preamble and judgement sections, includes diverse entity types detailed in their legal NER taxonomy, such as court, petitioner, respondent, and statute, among others. The training set, drawn from judgements between 1950 and 2017, features 29,964 entities and the development and test sets, spanning 2018 to 2022. This dataset not only facilitates training and evaluation of NER models specific to the legal domain but also provides a structured framework for assessing the performance of NER systems on legal texts. Their approach leverages a combination of manual annotation and ML techniques to ensure the precision of entity recognition in legal judgements.
10. Legal Argument Mining
Legal argument mining is the use of natural language processing to identify and extract arguments from legal documents. It aims to automate the extraction of argumentative content, thereby improving performance in legal research and practice. Current research focuses on developing domain-specific annotation schemes and employing advanced deep learning models to tackle the complexities of legal language and argument structures.
10.1. Task
Palau and Moens (2009) present pioneering research in LAM, focusing on detecting, classifying, and structuring arguments within legal texts. Their work introduces methods to automatically identify arguments, distinguish argumentative from non-argumentative sentences, and classify argumentative propositions as either premises or conclusions using statistical classifiers such as maximum entropy models, Naive Bayes classifiers, and SVM. The authors utilise the Araucaria and the ECHR corpora for evaluation. Results demonstrate that statistical methods can distinguish argumentative sentences, achieving approximately 80% accuracy on the ECHR corpus. Additionally, they discuss methods to resolve argument segmentation challenges through structural and semantic analyses, proposing that semantic relatedness measures (based on ontology or corpus-derived statistics) can enhance argument boundary detection. Finally, they investigate detecting argument structures through rhetorical pattern analysis and suggest employing context-free grammars as an initial step toward full argumentative parsing.
Zhang et al. (2023b) propose a graph-based framework for LAM that replaces the traditional pipeline approach with an end-to-end architecture. By modelling each legal document as a graph—where nodes represent text segments and edges capture sequential or semantic relationships—they mitigate error propagation across sub-tasks. Their method employs virtual node graph augmentation, which adds a global node connected to all text segments, and a collective classification algorithm that iteratively refines predictions using neighbouring node labels. They evaluate on the ECHR and the Court of Justice of the European Union (CJEU) datasets using Graph Convolutional Network (GCN) and Residual Gated Graph ConvNet (ResGCN) models. In particular, ResGCN demonstrates better performance on both datasets, surpassing baseline methods in classifying premises, conclusions, and non-argumentative text.
Santin et al. (2023) introduce a novel approach for predicting argumentation structures in CJEU decisions on fiscal state aid. They propose a tailored model, ResAttArg, which integrates residual connections, attention mechanisms, and multi-task learning to jointly predict argument component types and their inferential links. The model leverages domain-specific features—such as pre-trained word embeddings and encoding of inter-component distances—to capture complex argumentative relations. In contrast to a fine-tuned transformer baseline, the ensemble of ResAttArg networks shows improvements in linking argumentative components. The authors also observe that conventional oversampling strategies can degrade performance and that expanding the candidate span does not yield additional benefits. In addition, this study expand their previous dataset, Demosthenes, which consists of 40 decisions from CJEU on fiscal state aid. The corpus is annotated with detailed argumentative components—such as premises and conclusions—and is enriched with an annotation layer that captures inferential links and relation subtypes.
Habernal et al. (2024) utilise advanced NLP techniques to mine legal arguments from ECHR decisions. Specifically, they employ a multi-task transformer-based model, which is a deep learning approach designed for both argument identification and classification. This model leverages the capabilities of transformers—a widely used architecture in modern NLP—to process and understand complex legal texts, outperforming previous legal NLP models according to expert evaluations.
10.2. Datasets
Poudyal et al. (2020) present the ECHR corpus for LAM, a structured dataset of 42 decisions from ECHR, annotated with argumentative components: premises, conclusions, and non-argumentative text. The corpus facilitates research on argument mining by enabling three key tasks: argument clause recognition, clause relation prediction, and premise/conclusion classification. The annotation process involved legal experts, with iterative refinements leading to 80% inter-annotator agreement. Habernal et al. (2024) work on the same dataset and expand it to 373 decisions related to Articles 3, 7, and 8 of the ECHR, totalling 2.3 million tokens with 15,205 annotated argument spans. Annotated by six law students on the INCepTION [^8] https://inception-project.github.io platform under expert supervision, it achieved high reliability (Krippendorff’s alpha $\approx$ 0.80).
11. Large Legal Corpora
Training LLMs for legal NLP requires extensive legal corpora that are transparent in sourcing, safeguard privacy, and minimise bias to ensure fairness and accuracy. These corpora serve as the foundation for developing models capable of handling diverse legal tasks. To assess model performance, evaluation benchmarks provide structured datasets and standardised metrics for tasks such as judgement prediction, QA, case retrieval, and entailment. Together, large legal corpora and evaluation benchmarks try to support the advancement of reliable legal AI applications.
11.1. Evaluation benchmarks
@@
Zheng et al. (2021) introduce the CaseHOLD dataset, a novel benchmark for evaluating NLP models in the legal domain, designed to address the challenge of identifying the legal holdings from case texts. The dataset contains over 53,000 multiple choice questions derived from U.S. case law citations, where each question requires the identification of the correct holding from a set of potential answers. This task, simulating a fundamental lawyering skill taught in law school, involves contextual understanding and application of legal rules to factual situations. CaseHOLD is aimed at enhancing model training by focusing on semantic matching and the ability to discern legal principles. The dataset is structured to provide a challenging yet accessible resource for NLP researchers, with a clear focus on promoting deeper understanding and application of legal rules in automated systems. It utilises a format where a cited text serves as a prompt with five answer options—one correct holding and four closely related incorrect holdings—to refine the models’ abilities to reflect legal reasoning.
@@
Chalkidis et al. (2022a) introduce the Legal General Language Understanding Evaluation (LexGLUE) benchmark, a comprehensive suite of datasets aimed at assessing the capabilities of NLP models across various legal tasks. The benchmark covers datasets, such as ECtHR (Chalkidis et al., 2019a), SCOTUS [^9] https://www.supremecourt.gov, EUR-LEX, LEDGAR (Tuggener et al., 2020), UNFAIR-ToS (Lippi et al., 2019), and CaseHOLD (Zheng et al., 2021), each chosen for its complexity, relevance, and need for legal expertise. These datasets cover a range of tasks from multi-label and multi-class classification to multiple-choice questions and are split chronologically into training, development, and test sets to provide standardised evaluation metrics. For instance, ECtHR datasets focus on violations of European Convention of Human Rights provisions, SCOTUS database classifies U.S. Supreme Court opinions by legal issues, EUR-LEX database involves labelling EU laws with EuroVoc concepts, LEDGAR classifies provisions of U.S. contracts, UNFAIR-ToS identifies unfair terms in online service agreements, and CaseHOLD involves QA about legal rulings.
@@
Chalkidis et al. (2022b) present FairLex, a benchmark suite consisting of four legal datasets—ECtHR (Chalkidis et al., 2019a), SCOTUS, FSCS, and CAIL—that address the fairness of NLP applications across diverse legal jurisdictions and languages, including English, German, French, Italian, and Chinese. These datasets, curated from European Council, USA, Switzerland, and China, cover various legal tasks such as LJP, issue area classification, and crime severity prediction, aiming to test the performance and fairness of LMs in recognising and classifying legal texts. FairLex focuses on ensuring demographic, regional, and legal topic fairness by analysing attributes, such as gender, age, region of origin, and legal areas within cases. Each dataset in FairLex provides a substantial number of cases, systematically divided into training, development, and test sets, and includes detailed attributes, such as the defendant state in ECtHR, decision direction in SCOTUS, legal areas in FSCS, and demographic details in CAIL.
@@
Rabelo et al. (2022) summarise the 8th Competition on Legal Information Extraction and Entailment (COLIEE 2021), which featured five tasks across case and statute law, engaging participants from various teams to apply diverse NLP approaches. The competition tasks included case law retrieval and entailment, as well as statute law retrieval and entailment with and without prior retrieved data. Specifically, task 1 focused on extracting relevant supporting cases from a corpus, while task 2 involved identifying paragraphs from cases that entail a given new case fragment. For statute law, tasks 3 and 4 entailed retrieving and answering questions based on civil code statutes, with task 5 challenging participants to answer without pre-retrieved statutes. The datasets used varied in complexity, from 4415 case files in task 1 with a need to identify noticed cases without relying on citations, to the civil code-based tasks 3, 4, and 5 which adapted to recent legal revisions in Japanese law and excluded untranslated parts, reflecting the ongoing evolution and challenge in legal NLP applications.
@@
Barale et al. (2023) present AsyLex, a pioneering dataset tailored for refugee law applications, featuring 59,112 documents from Canadian refugee status determinations spanning from 1996 to 2022. This dataset is designed to enhance the capabilities of NLP models in legal research by providing 19,115 gold-standard human-annotated and 30,944 inferred labels for entity extraction and LJP. Key contributions include anonymizing decision documents, employing a robust annotation methodology, and creating datasets for specific NLP tasks, such as entity extraction and judgement prediction. This rich corpus, with detailed annotations across 22 categories, supports complex legal NLP tasks, thereby filling the gap in resources for the legal domain.
@@
Niklaus et al. (2023) introduce LEXTREME, a multilingual benchmark specifically designed to evaluate LMs on legal NLP tasks, a critical step given the unique challenges of legal language. Surveying legal NLP literature from 2010 to 2022, they curate 11 datasets spanning 24 languages and cover a variety of legal domains, employing datasets that only involve human-annotated texts or those with annotations derived through clear methodological frameworks. They introduce two aggregate scores to facilitate fair comparison across models: the dataset aggregate score and the language aggregate score, revealing a performance correlation with model size on LEXTREME. The benchmark consists of three task types: single label text classification, multi label text classification, and NER, using existing splits for training, validation, and testing when available, or creating random splits otherwise. This effort marks an advancement in testing NLP capabilities across a diverse range of legal documents and languages.
@@
Park and James (2023) explore the creation of a Natural Language Inference dataset within the legal domain, focusing on criminal court verdicts in Korean. Their methodology includes the innovative use of adversarial hypothesis generation to challenge annotators and enhance the robustness of the dataset, supported by visual tools for hypothesis network construction. The data collection involves extracting context from verdicts and augmenting it using Easy Data Augmentation (Wei and Zou, 2019) techniques and round-trip translation to generate a dataset for training and testing Natural Language Inference models. The study highlights issues such as annotators’ limited domain knowledge and challenges in handling long contexts but provides solutions, such as targeted data collection and the use of gamification to boost annotator engagement and productivity.
@@
Goebel et al. (2024) summarise COLIEE 2023, featuring four tasks across case and statute law with participation from ten different teams engaging in multiple tasks. task 1 involves legal case retrieval, requiring participants to extract supporting cases from a corpus, and task 2 focuses on legal case entailment, identifying paragraphs that entail aspects of a new case. Task 3 and 4, based on Japanese civil code statutes from the bar exam, involve retrieving relevant articles and verifying statements, respectively. The competition leverages a dataset of over 5,700 case law files and introduces new query cases and test questions sourced from recent bar exams, testing the efficacy of different teams’ approaches in handling complex legal texts and hypotheses in a controlled competitive environment.
@@
Östling et al. (2024) introduce the Cambridge Law Corpus (CLC), a legal dataset featuring 258,146 cases from UK courts, dating from the 16th century to the present. The corpus includes raw text and metadata across various court types, and is structured in XML format for ease of use and annotated for case outcomes in a subset of 638 cases. Additionally, the CLC is supported by a Python library for data manipulation and ML applications.
11.2. Large corpora for pre-training
Henderson et al. (2022) introduce the ‘Pile of Law’, the first and an important large corpus in the legal domain, containing a 256 GB dataset of open-source English-language legal and administrative data. This dataset includes contracts, court opinions, legislative records, and administrative rules, curated to explore data sanitation norms across legal and administrative settings and serve as a tool for pre-training legal-domain LMs. They emphasise the legal norms governing privacy and toxicity filtering, detailing how the dataset reflects these norms through built-in filtering mechanisms in the collected data, which include court filings, legal analyses, and government publications. By analysing how legal and administrative entities handle sensitive information and potentially offensive content, the paper provides actionable insights for researchers to improve content filtering practices before pre-training LLMs, thereby enhancing the ethical use of NLP in legal applications.